### v3

# Carlo Orientale Caputo SISSA - International School for Advanced Studies, 34136 Trieste, Italy Elias Seiffert

Abstract
Abstraction is the process of extracting the essential features from raw data while ignoring irrelevant details. This is similar to the process of focusing on large-scale properties, systematically removing irrelevant small-scale details, implemented in the renormalisation group of statistical physics. This analogy is suggestive because the fixed points of the renormalisation group offer an ideal candidate of a truly abstract – i.e. data independent – representation.
It has been observed that abstraction emerges with depth in neural networks. Deep layers of neural network capture abstract characteristics of data, such as ”cat-ness” or ”dog-ness” in images, by combining the lower level features encoded in shallow layers (e.g. edges). Yet we argue that depth alone is not enough to develop truly abstract representations. We advocate that the level of abstraction crucially depends on how broad the training set is. We address the issue within a renormalisation group approach where a representation is expanded to encompass a broader set of data. We take the unique fixed point of this transformation — the Hierarchical Feature Model – as a candidate for an abstract representation. This theoretical picture is tested in numerical experiments based on Deep Belief Networks trained on data of different breadth. These show that representations in deep layers of neural networks approach the Hierarchical Feature Model as the data gets broader, in agreement with theoretical predictions.
Following Marr [1], one may argue that the conceptual underpinnings of cognitive functions are independent of whether they are implemented in-silico or in a biological brain. In this spirit, this paper adopts artificial neural networks (ANN) as a playground to explore how abstraction arises from the process of making sense of raw complex data. To illustrate this, consider vision. Both in biological brains [1] and in ANNs [2], vision involves a hierarchical organization of representations: shallow layers detect low-level features, such as edges [3], while deeper layers integrate these features to recognize more abstract, higher-order constructs like objects and faces [2]. In particular, deeper layers are capable of recognising an object or a face irrespective of it’s position, orientation, scale, or of context [^1] This ability can be promoted in ANN by either augmenting the data using invariances [4] or by explicitly implementing them in their architecture – as in convolutional neural networks [2]. Yet even simple neural networks are able to develop a convolutional structure by themselves [5]. [6, 7]. This parallel underscores the broader principle that abstraction emerges through layered processing, regardless of the underlying substrate on which the process is implemented.
In general, abstract representations extracted from data have been discussed in terms of “cognitive maps” [8, 9]. A cognitive map is not only an efficient and flexible scaffold of data, but it is also endowed with a structure of relations – uncovered from the data – that enables abstract computation [^2] Relational structures such as “Alice is the daughter of Jim” and ”Bob is Alice’s brother” allow for computations (e.g. ”Jim is Bob’s father”) which are invariant with respect to the context (Alice, Jim and Bob can be replaced by any triplet of persons that stand in the same relation, in this example) [8]. [8, 9] and supports complex functions. For example, spatial navigation relies on the representation built by several assemblies of specialised neurons, such as grid cells [10, 11].
At even higher levels of cognition, representations should integrate data from a broader set of domains, or perceptual modalities [12], each of which may be organised according to cognitive maps of a different nature. For example, while visual stimuli are described by object manifolds with supposedly local euclidean topology [13, 14], odours have been suggested to be organised in hyperbolic spaces [15]. Higher order representations that integrate the two should therefore be even more abstract, i.e. independent of the data.
A lot has been understood about the role of depth in learning [^3] Strictly speaking, most of these insights pertain to supervised learning, yet we assume they reveal properties that are relevant also for unsupervised learning, which is our focus. [13, 16, 17, 14, 18, 19, 20]. In particular, depth exploits the compositional structure of the data boosting training performance [16, 18, 20] and classification capacity [17]. Indeed, inner layers of ANN portray data that correspond to the same object as “object manifolds” [13] that become better and better separable with depth [17], while extracting hierarchies of features that promote taxonomic abstraction [^4] Taxonomic abstraction is based on the idea that objects are similar if they share the same features (e.g. ”cats” and ”dogs” both have four legs, a tail, etc) and belong to the same category (mammals). Taxonomic abstraction is fundamentally distinct from thematic abstraction, which is based on co-occurrence between objects that share no features (e.g. ”cat” and ”sofa”), as discussed in [21]. In deep neural networks [18, 20] shallow layers capture statistical associations (thematic) while deep layers encode compositional, taxonomic structures. A similar transition is observed in humans with development: while children favour thematic abstractions, adults more frequently rely on taxonomic (or categorical) structures [22]..
A possible road to abstraction has been suggested [23, 24] exploiting the analogy between the processing of data in deeper and deeper layers and the renormalisation group (RG) in statistical physics [25] (see Fig. 1 A). This analogy is based on the observation that higher order features in learning are akin to large scale properties in statistical models, which are those that the RG singles out. This idea is suggestive for at least two reasons: First the RG is the theoretical tool to study critical phenomena [25] and both artificial and biological learning exhibit critical features [26, 27, 28, 29, 30, 31]. Second, and most importantly, repeated RG transformations lead to universal distributions, which are an ideal candidate for abstract representations.
In this paper we argue that this analogy misses a fundamental ingredient, which is breadth, i.e. the diversity in the input data or stimuli [^5] We distinguish breadth from width, a term commonly used in the literature to denote the number of variables in different layers.. It is common sense that a representation encompassing a broader universe of circumstances should be more abstract than one describing only a limited domain. Likewise, we argue that representations in higher levels of the cognitive hierarchy, that integrate a broader set of stimuli, should be more abstract, i.e. independent of the data. This paper approaches the problem of characterising such abstract representations on the basis of their sole statistical properties, with no reference to what is being represented. In this respect, we depart from the typical “tuning curve” approach in computational neuroscience, in which levels of abstraction are assessed in terms of the features of the data – e.g. edges or faces – that a representation encodes, as well as from other efforts aimed at finding those properties that make structured data learnable (see e.g. [18, 20, 32]). Focusing on the marginal distribution of activations of inner layers allows us to characterise abstract representations purely through information-theoretic principles. We focus on the simple case of static data: a representation in this paper is a probability distribution over a set of binary variables.
Figure 1: Illustrative example of the RG in statistical physics (A) and of it’s application in learning (B and C). The RG (A) entails a coarse gaining (or decimation) step, whereby low scale degrees of freedom are integrated out, e.g. with the introduction of block variables (large dots in the middle A-panel), and a rescaling step that restores the original size of the system. In a representation of a given domain of items (animals), in B coarse graining is performed zooming into those with a specific feature (living in water) and rescaling corresponds to enriching the representation by adding further details. The same procedure can be reversed (in C): the representation describing a particular domain (animals from planet Earth) is retrained on a wider domain (animals from many planets), neglecting small scale details (e.g. the difference between whales and dolphins).
In this setting, we show that abstract representations emerge as the fixed point of a RG transformation whereby a representation for a given domain is updated to describe data from a broader domain. In this process, assuming a constant coding cost, low level details are sacrificed in order to make space for high level features describing the organisation of the data within the broader domain. Fig. 1 C sketches this process for an illustrative example. This process of zooming out to a broader domain while loosing low level details can also be inverted, by zooming into a specific part of the data, thus uncovering low level details (see Fig. 1 B). We show that in both cases, the transformation has a unique fixed point that coincides with the Hierarchical Feature Model (HFM), recently introduced in [33]. This is reassuring for at least two reasons: First the HFM is a maximum entropy model fully determined by a single sufficient statistics, which is the average level of detail of the features, or the coding cost. This is indeed the only relevant variable in an abstract representation. Second, the HFM satisfies the principle of maximal relevance. The relevance has been recently introduced [34] as a quantitative measure of “meaning” that captures Barlow’s intuition [35] that meaning is carried by redundancy. We refer to Section 1.1 for a brief discussion of the relevance, or to Ref. [36] for an extended account. Let it suffice to say that internal representations of well trained learning machines have been shown to satisfy the principle of maximal relevance [37, 38].
While a transformation in terms of breadth has a unique fixed point, we shall also argue that a similar transformation in terms of depth alone may have many fixed points. This is an important observation for the second part of the paper, where we shall search for evidence of our theory in empirical data of deep belief networks trained on different datasets. Our numerical experiments, presented in Section 3, will show that convergence to the abstract representation of the HFM requires the combined effect of depth and breadth. Section 4 discusses the results and provides some concluding remarks. All technical details are relegated to the Appendix [^6] The present paper supersedes the preliminary results presented in the Master thesis of one of us [39]..
1 The framework
In this paper, a representation is a probability distribution $p(\mathbf{s})$ over a string of binary variables $\mathbf{s}=(s_{1},\ldots,s_{n})$ ($s_{i}\in\{0,1\}$), that model the activation of “neurons” inside a learning machine. For example, $\mathbf{s}=\mathbf{s}^{(\ell)}$ can be the variables in the $\ell^{\rm th}$ layer of a neural network with many layers. We shall focus on unsupervised learning and generative models such as a Deep Belief Network (DBN), that mathematically can be seen as a joint probability distribution
$p(\mathbf{x},\mathbf{s}^{(1)},\ldots,\mathbf{s}^{(L)})=p(\mathbf{x}|\mathbf{s}^{(1)})p(\mathbf{s}^{(1)}|\mathbf{s}^{(2)})\cdots p(\mathbf{s}^{(L-1)}|\mathbf{s}^{(L)})p(\mathbf{s}^{(L)})$ (1)
of the variables $\mathbf{x}=(x_{1},\ldots x_{m})$ in the visible layer and the variables $\mathbf{s}^{(\ell)}=(s_{1},\ldots,s_{n_{\ell}})$ in the $L$ hidden layers ($\ell=1,\ldots,L$). Eq. (1) also highlights the Markovian nature of information processing whereby the activation of layer $\ell$ only depends on data through layer $\ell-1$. We focus on this probability distribution after the network has been trained on some data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ by maximising the log-likelihood $\mathcal{L}(\hat{\mathbf{x}})=\sum_{i}\log p(\mathbf{x}_{i})$ over the parameters of the network. We refer to Appendix A for a more detailed discussion of such networks. For our purposes, let it suffice to say that the object of our study will be the marginal distribution $p(\mathbf{s}^{(\ell)})$ of $\mathbf{s}^{(\ell)}$ after training. In a well trained network, sampling $\mathbf{s}^{(\ell)}$ from this distribution and propagating the state until the visible layer generates data points which are statistically indistinguishable from the data.
We may think of $s_{i}^{(\ell)}$ as indicator variables of abstract features: i.e. $s_{i}^{(\ell)}=1$ generates points with $-\ell$ feature $i$ and $s_{i}^{(\ell)}=0$ does not. In what follows, we shall drop the index $\ell$ of the layer when not needed. We shall also use, when needed, the notation $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ to keep track of indices and $\mathbf{0}$ (or $\mathbf{0}_{1:n}$) to indicate the featureless state, i.e. the one with $s_{i}=0$ for all $i=1,\ldots,n$, which, as we shall see, describes most common objects.
In order to set the stage, we shall first recall few notion on the relevance and on the Hierarchical Feature Model.
1.1 The relevance
We describe representations $p(\mathbf{s})$ in terms of the variable $E_{\mathbf{s}}=-\log_{2}p(\mathbf{s})$, which is the minimal number of bits needed to represent state $\mathbf{s}$. The average coding cost $H[\mathbf{s}]=\mathbb{E}\left[E_{\mathbf{s}}\right]$ is the usual Shannon entropy and counts the number of bits needed to describe one point of the dataset. Following Ref. [36], we shall call $H[\mathbf{s}]$ the resolution.
The resolution $H[\mathbf{s}]$ is a measure of information content but not of information “quality”. Meaningful information should bear statistical signatures that allow it to be distinguished from noise. These make it possible to identify relevant information before finding out what that information is relevant for, a key feature of learning in living systems. Following Ref. [36], we take the view that the hallmark of meaningful information is a broad distribution of coding costs. Here breadth can be quantified by the relevance, which is the entropy of the coding cost $E_{\mathbf{s}}$
$H[E]=-\sum_{E}p(E)\log_{2}p(E)\,,$ (2)
where $p(E)=W(E)2^{-E}$ is the probability that a state $\mathbf{s}$ randomly drawn from $p(\mathbf{s})$ has $E_{\mathbf{s}}=E$, and $W(E)$ is the number of states $\mathbf{s}$ with $E_{\mathbf{s}}=E$. The principle of maximal relevance [36] postulates that maximally informative representations should achieve a maximal value of $H[E]$, which correspond to a uniform distribution of coding costs ($p(E)={\rm const}$ or $W(E)=W_{0}2^{E}$). Representations where coding costs are distributed uniformly should be promoted for the reason that, in an optimal representation, the number $W(E)$ of states $\mathbf{s}$ that require $E$ bits to be represented should match as closely as possible the number ($2^{E}$) of codewords that can be described by $E$ bits. Representation of maximal relevance with a given resolution also have an exponential degeneracy of states $W(E)=W_{0}e^{\nu E}$, with $\nu$ that depends on $H[\mathbf{s}]$. Note that states $\mathbf{s}$ and $\mathbf{s}^{\prime}$ with very different coding costs $E_{\mathbf{s}}$ and $E_{\mathbf{s}^{\prime}}$ can be distinguished by their statistics, because they would naturally belong to different typical sets [^7] By the law of large numbers, typical samples of weakly interacting variables all have approximately the same coding cost, a fact knowns as the asymptotic equipartition property [40]. Following Ref. [41], we take the view that a trained DBN distinguishes the points in a dataset in different typical sets.. Representations that maximise the relevance harvest this benefit in discrimination ability that is accorded by statistics alone.
1.2 The Hierarchical Feature Model
The HFM encodes the principle of maximal relevance. It describes the distribution of a string $\mathbf{s}_{1:n}=(s_{1},\ldots,s_{n})$ of binary variables that we take as indicators of whether each of $n$ features is present ($s_{i}=1$) or not ($s_{i}=0$). Features are organised in a hierarchical scale of detail and we require that the occurrence of a feature $s_{k}=1$ at level $k$ does not provide any information on whether lower order features are present or not. This means that, conditional on $s_{k}=1$, all lower order features are as random as possible, i.e. $H[\mathbf{s}_{1:k-1}|s_{k}=1]=k-1$ in bits. This requirement implies that the Hamiltonian $E_{\mathbf{s}}$ should be a function of $m_{\mathbf{s}}=\max\{k:~{}s_{k}=1\}$, with $m_{\mathbf{s}}=0$ if $\mathbf{s}=\mathbf{0}$ is the featureless state [^8] This is because $p(\mathbf{s}|m_{\mathbf{s}}=k)=2^{-k+1}$ so that $p(\mathbf{s})=p(\mathbf{s}|m_{\mathbf{s}})p(m_{\mathbf{s}})=p(m_{\mathbf{s}})/2^{m_{\mathbf{s}}-1}$. ($s_{i}=0~{}\forall i$). Since there are $2^{k-1}$ states with $m_{\mathbf{s}}=k$, the principle of maximal relevance (i.e. the requirement that $W(E)=W_{0}e^{\nu E}$) excludes all functional forms between $E_{\mathbf{s}}$ and $m_{\mathbf{s}}$ that are not linear. This leads to the HFM, that assigns a probability
$h_{n}(\mathbf{s})=\frac{1}{Z_{n}}e^{-gm_{\mathbf{s}}}\,,$ (3)
to state $\mathbf{s}$, where the partition function $Z_{n}$ ensures normalisation. We refer to [33] for a detailed discussion of the properties of the HFM. In brief, in the limit $n\to\infty$ the HFM features a phase transition at $g_{c}=\log 2$ between a random phase where $H[\mathbf{s}]$ is of order $n$ for $g<g_{c}$, and a “low temperature” phase where $h_{n}(\mathbf{s})$ is dominated by a finite number of states (and $H[\mathbf{s}]$ is finite in the limit $n\to\infty$).
Marginalising over the low order features $\mathbf{s}_{1:k}$ returns a mixture between a HFM over the remaining $n-k$ features and a frozen state
$\sum_{\mathbf{s}_{1:k}}h_{n}(\mathbf{s}_{1:n})=\frac{\xi^{k}Z_{n-k}}{Z_{n}}h_{n-k}(\mathbf{s}_{k+1:n})+\frac{(1-\xi/2)(\xi^{k}-1)}{(\xi-1)Z_{n}}\delta_{\mathbf{s}_{k+1:n},\mathbf{0}_{k+1:n}}\,.$ (4)
On the other hand, marginalising over the high order ones yields a mixture between the HFM and the uniform (maximum entropy) distribution
$\sum_{\mathbf{s}_{k+1:n}}h_{n}(\mathbf{s}_{1:n})=\frac{Z_{k}}{Z_{n}}h_{k}(\mathbf{s}_{1:k})+\left(1-\frac{Z_{k}}{Z_{n}}\right)2^{-k}\,.$ (5)
2 Renormalization transformations
The renormalisation group (RG) [25] translates the process of focusing on large scale properties of statistical models into a mathematical formalism. As shown in Fig 1 A), this process is typically composed of two parts: i) coarse graining, whereby small scale details are eliminated and ii) rescaling in order to restore the original (length) scale. The combined effect of these two steps is a transformation from a probability distribution $p(\mathbf{s})$ to a different one $p^{\prime}=\Re[p]$ whose fixed points $p^{*}=\Re[p^{*}]$ describe scale invariant states. In statistical physics, fixed points are endowed with universal properties which make them depend solely on few fundamental characteristics, such as symmetries and conservation laws, space dimension and dimensionality of the relevant variables (order parameters) [25]. In the context of learning, such universal distributions are natural candidates for an abstract representation, and the RG offers the ideal conceptual framework to search for them.
In this section we discuss the application of these ideas with respect to depth and then to breadth.
2.1 The limit of infinite depth
The similarity between the extraction of higher order features in deeper layer of neural networks and the RG [25] in statistical physics has been pointed out since long [23]. Indeed, Koch et al. [24] argue that training DBNs over configurations of an Ising model generates a distribution in a hidden layer which is similar to that obtained in the coarse graining step of the RG applied to the distribution of the previous layer. Yet the transformation of $p(\mathbf{s})$ from one layer to the next that is implemented in training does not account for the rescaling step. Typically, training compresses the representation, i.e. the number $n_{\ell}$ of variables and the entropy $H[\mathbf{s}^{(\ell)}]$ of layer $\ell$ are smaller than those of layer $\ell-1$. One could envisage different ways to restore the values of $n_{\ell-1}$ and $H[\mathbf{s}^{(\ell-1)}]$, but their interpretation in terms of learning is not that transparent [^9] One way to expand the number of variables while preserving information content is to introduce parity checks. This leaves $H[\mathbf{s}^{(\ell)}]$ invariant. The entropy could be restored to its value in the previous layer by a large deviation transformation $p(\mathbf{s})\to Ap^{\beta}(\mathbf{s})$ that is analogous to a change in temperature, or by introducing randomness with e.g. a binary symmetric channel [40]. We have explored numerically the combination of these two steps in order to implement rescaling, but we failed to find a stable fixed point. Note, in particular, that the introduction of parity checks injects high order statistical dependencies in $p(\mathbf{s})$ which are generally hard to detect in training..
This Section argues that the transformation across many layers does not lead to universal distributions. Rather, we argue that a representation converges to one of many sharply localised absorbing states in this process. Here we provide a simple argument: consider a bounded function $\phi_{0}(\mathbf{x})\in[\phi_{-},\phi_{+}]$ of the data. The transformation
$\phi_{\ell}(\mathbf{s}^{(\ell)})=\sum_{\mathbf{x}}\phi_{0}(\mathbf{x})p(\mathbf{x}|\mathbf{s}^{(\ell)})$ (6)
and the distribution Eq. (1) defines a sequence of random variables $\phi_{\ell}$, $\ell=0,1,\ldots,L$, which is a martingale, i.e.
$\mathbb{E}\left[\phi_{\ell}(\mathbf{s}^{(\ell)})|\phi_{\ell-1}(\mathbf{s}^{(\ell-1)})=\phi_{\ell-1}\right]=\phi_{\ell-1}\,.$ (7)
[…]
In summary, our analysis suggests that rather than a single fixed point, depth promotes the emergence of a representation characterised by a mixture of sharply peaked fixed points, each of which describes a subset of the training data.
2.2 The limit of infinite breadth
The internal representation $p(\mathbf{s})$ of an internal layer of a learning machine depends on its depth but also on the data $\hat{\mathbf{x}}$ used in training. We focus on how a representation changes when it is trained on a larger dataset $\hat{\mathbf{x}^{\prime}}$ incorporating data from a broader domain (Fig. 1 C) or when the network is trained only on a subset of the data (Fig. 1 B). An illustrative example, that we shall discuss later in the empirical part, is that of networks trained on images of handwritten digits, or of all characters (including letters), or of only one digits (e.g. $2$).
2.2.1 Coarse graining
We describe a transformation whereby a representation defined in terms of $n$ hierarchical features is transformed zooming out to include a further coarse grained feature. This transformation describes how the internal representation $p(\mathbf{s})$ of a deep layer changes when the universe of objects it represents is expanded to include further objects. Such an expansion is analogous to the rescaling step in the RG (the rightmost in Fig. 1 A). In the RG the configuration of the larger system is drawn from the Boltzmann distribution with the rescaled couplings, which is a maximum entropy distribution. Likewise, when the universe of objects is expanded we rely on a maximum ignorance assumption on how the new coarse grained feature is distributed. This reflects the principle that learning should venture into the unknown with no prejudice.
At the same time, a limited information capacity imposes to discard fine details in the description provided by the representation marginalising over the most fine grained feature. This step is also accompanied by a redefinition of the palette of features so as to restore the original value $H[\mathbf{s}]$ of the resolution. This corresponds to the coarse graining step in the RG (the leftmost in Fig. 1 A).
This transformation can be seen as combining both depth and breadth in the coarse graining and rescaling steps, respectively. The coarse graining step involves training the internal representation $\mathcal{R}=\mathcal{T}(\mathcal{D})$ of a learning machine on a dataset $\mathcal{D}$. This transformation $\mathcal{D}\Rightarrow\mathcal{R}$ corresponds to a change in depth and it returns a compressed representation $\mathcal{R}$ of the data $\mathcal{D}$. In order to restore the original dimensions of the problem, we add a rescaling step where the data $\mathcal{D}\Rightarrow\mathcal{D}\cup\mathcal{D}^{\prime}$ is expanded to encompass a wider domain, in such a way that the representation $\mathcal{R}^{\prime}=\mathcal{T}(\mathcal{D}\cup\mathcal{D}^{\prime})$ learned from the expanded dataset has the same dimension of the original data $\mathcal{D}$. The representation $\mathcal{R}^{\prime}$ generates the data $\mathcal{D}$ for the next step of the RG procedure.
As suggested in [33], such an expansion in the breadth of data learned by the representation comes with a substantial redefinition of features in a shallow network. Conversely, we assume that in deep layers of a neural network this transformation involves smooth changes of intermediate features. The existence of a fixed point then ensures continuity of $p(\mathbf{s})$ in this process. In other words, while what each feature represents in the data, i.e. $p(\mathbf{x}|\mathbf{s})$ may change considerably, the way in which features are organised, i.e. $p(\mathbf{s})$, does not. Ref. [33] argues that such a continuity property provides advantages which make learning more similar to understanding.
Let $p(\mathbf{s}_{1:n})$ be a representation. We envisage a transformation $p\to p^{\prime}=\Re_{\uparrow}[p]$ based on the following steps:
  1. 1.
    Marginalise $s_{n}$
    $p(\mathbf{s}_{1:n-1})=\sum_{s_{n}=0,1}p(\mathbf{s}_{1:n})\,.$ (8)
    In this step, the most detailed feature is eliminated, analogously to what happens when processing data from one layer to next in a deep neural network.
  2. 2.
    Add a new random feature with $p(s_{0}=1)=p(s_{0}=0)=\frac{1}{2}$, i.e.
    $\tilde{p}(\mathbf{s}_{0:n-1})=\frac{1}{2}p(\mathbf{s}_{1:n-1})$ (9)
    In this step, the representation is expanded to describe a wider universe of objects. The distribution of the new feature is independent of $\mathbf{s}_{1:n-1}$. This captures a genuine discovery process characterised by a large scale organisation (described by $s_{0}$) which cannot be described in terms of combinations of already known features.
  3. 3.
    Shift indices $\mathbf{s}_{1:n}^{\prime}=\mathbf{s}_{0:n-1}$
  4. 4.
    Renormalise
    $p^{\prime}(\mathbf{s}_{1:n}^{\prime})=(1-\alpha)\tilde{p}(\mathbf{s}_{0:n-1})+\alpha\delta_{\mathbf{s}_{1:n}^{\prime},\mathbf{0}_{1:n}}$ (10)
    where $\alpha$ should be fixed so that the coding cost $H[\mathbf{s}_{1:n}]=H[\mathbf{s}_{1,n}^{\prime}]$ remains the same. Notice that this last step implies a redefinition of typical objects, those described by the featureless state $\mathbf{s}=\mathbf{0}$.
In the first two steps the resolution $H[\mathbf{s}]$ increases, because the $n^{\rm th}$ feature is replaced with a totally random one. In the last step, the resolution decreases by mixing the $\tilde{p}$ with a constant $\mathbf{0}_{1:n}$. Hence there is a unique solution for $\alpha$ (see Appendix B for more details).
A simple argument shows that the transformation $p\to p^{\prime}=\Re_{\uparrow}[p]$ converges to a unique fixed point. This is because, as shown above, there is a monotonous relation between $H[\mathbf{s}]$ and $\alpha$. For a fixed $\alpha$, the transformation described above is linear, which means that it can be expressed as
$p^{\prime}(\mathbf{s}_{1:n}^{\prime})=\sum_{\mathbf{s}_{1:n}}p(\mathbf{s}_{1:n})T_{\mathbf{s}_{1:n},\mathbf{s}_{1:n}^{\prime}}$ (11)
where $\hat{T}$ is a stochastic matrix. The associated Markov chain describes a random walk with resetting [43] on the de Bruijn graph [44], and it is shown in Fig. 2 for $n=3$.
Figure 2: Graphical representation of the transition matrix $T_{\mathbf{s}_{1:n},\mathbf{s}_{1:n}^{\prime}}$ of the coarse graining RG for $n=3$. States are represented by circles and transitions by arrows. From each state $\mathbf{s}_{1:n}$ there are only three non-zero matrix elements of $T_{\mathbf{s}_{1:n},\mathbf{s}_{1:n}^{\prime}}$. Two of them correspond to the possible states $\mathbf{s}_{1:n}^{\prime}=(s_{0},\mathbf{s}_{1:n-1})$ that can be reached by either adding $s_{0}=1$ to the left of $\mathbf{s}_{1:n}$ (red links) or adding $s_{0}=0$ (blue dashed links). Both transitions occur with probability $T_{\mathbf{s}_{1:n},\mathbf{s}_{1:n}^{\prime}}=\frac{1-\alpha}{2}$. The third transition resets to the $\mathbf{0}_{1:n}$ state (green dotted links) and it occurs with probability $T_{\mathbf{s}_{1:n},\mathbf{0}_{1:n}}=\alpha$.
This Markov chain is clearly ergodic, because $\hat{T}^{m}$ has all strictly positive elements for $m\geq n$. This is because each time the variable $s_{0}$ is generated at random, so after $m$ iterations, every state $\mathbf{s}_{1:n}^{\prime}$ can be generated. By the theory of Markov chains [^11] We recall the Perron-Frobenius theorem, which states that a matrix with all positive elements has a unique maximal eigenvalue whose corresponding eigenvector has all positive elements. For an ergodic stochastic matrix this eigenvalue is one., under successive applications of the transformation, the distribution converges to a fixed point $p^{*}$ from any initial distribution $p$, and the limit is unique. The same necessarily applies to the transformation where $H[\mathbf{s}]$ is held fixed.
The unique fixed point $p^{*}$ of the coarse graining transformation is the marginal distribution of the $n$ most coarse grained features of an HFM with infinite features:
$p^{*}(s_{1:n})=\lim_{m\to\infty}\sum_{s_{n+1:m}}h_{m}(s_{1:m})\,.$ (12)
The proof of this result is shown in Appendix C. Notice that, by Eq. (5),
$\sum_{s_{n+1:m}}h_{m}(s_{1:m})=\frac{Z_{n}}{Z_{m}}h_{n}(s_{1:n})+\left(1-\frac{Z_{n}}{Z_{m}}\right)2^{-n}\,.$ (13)
For $g\leq g_{c}$ the distribution $p^{*}$ converges to the uniform distribution, because $Z_{n}/Z_{m}\to 0$ as $m\to\infty$. Therefore, for a finite $H[\mathbf{s}]<n$ (in bits) the fixed point has $g>g_{c}$.
2.2.2 Fine graining
The inverse transformation to the one described above, is obtained zooming in the representation of objects with $s_{1}=1$. In this process the universe of objects described is reduced but further fine grained details are added to the representation. The same general arguments as those discussed above for $\Re_{\uparrow}$ apply. The fine graining transformation $p\to p^{\prime}=\Re_{\downarrow}[p]$ is based on the following steps.
  1. 1.
    Zooming in on objects with $s_{1}=1$
    $p(\mathbf{s}_{2:n})=p(s_{1}=1,\mathbf{s}_{2:n})$ (14)
  2. 2.
    Shift indices $\mathbf{s}_{1:n-1}^{\prime}=\mathbf{s}_{2:n}$ and $\tilde{p}(\mathbf{s}_{1:n-1}^{\prime})=p(\mathbf{s}_{2:n})$.
  3. 3.
    Add a new feature $s_{n}^{\prime}$
    $p^{\prime}(\mathbf{s}_{1:n}^{\prime})=p^{\prime}(\mathbf{s}_{1:n-1}^{\prime}|s_{n})p(s_{n})$ (15)
    where $p(s_{n}=1)=q=1-p(s_{n}=0)$ sets the value of the entropy $H[\mathbf{s}_{1:n}^{\prime}]=H[\mathbf{s}_{1:n}]$ and we assume
    $p^{\prime}(\mathbf{s}_{1:n-1}^{\prime}|s_{n}=0)$ $=$ $\tilde{p}(\mathbf{s}_{1:n-1}^{\prime})$ (16)
    $p^{\prime}(\mathbf{s}_{1:n-1}^{\prime}|s_{n}=1)$ $=$ $2^{-n+1}\,.$ (17)
    The first of these two equations implies that the representation without the new feature is the same as the original representation over $\mathbf{s}_{2:n}$. The second equation enforces a maximum ignorance principle whereby the presence of the $n^{\rm th}$ feature ($s_{n}=1$) does not provide any information on whether more coarse grained features are present or not. This is the equivalent of the second step of the coarse graining procedure of Sect. 2.2.1, which assumes that higher level features do not impose constraints on lower level ones [^12] Indeed, by Eq. (9), $p^{\prime}(s_{1}^{\prime}|\mathbf{s}_{2:n}^{\prime}\neq\mathbf{0}_{2:n})=p(s_{1}^{\prime})=\frac{1}{2}$..
We can appeal to the same arguments as in Section 2.2.1 to show that the transformation $p\to p^{\prime}=\Re_{\downarrow}[p]$ has a unique fixed point. It is also easy to check that the fixed point is the HFM, with a value of $g$ that depends on $H[\mathbf{s}]$ (or $q$). Indeed $h_{n}(s_{1}=1,\mathbf{s}_{2:n})=h_{n-1}(\mathbf{s}_{2:n})$ and the HFM satisfies the condition (17) by definition.
3 Empirical evidence in Deep Belief Networks
In this Section we test the ideas discussed in previous Sections on Deep Belief Networks (DBNs) trained on different datasets.
3.1 Datasets, architectures and statistics
All our numerical experiments were run on variants of the MNIST dataset of handwritten digits, that we refer to with the letter M. Dataset 2-M was obtained from the data points of MNIST which correspond to the digit $2$, augmented by symmetry transformations, so as to have $=6\times 10^{4}$ data points. Dataset EM corresponds to the extended MNIST dataset, which combines digits and letters. We refer to Appendix A for more details on the datasets.
We trained Deep Belief Networks (DBNs) by successively training the Restricted Boltzmann Machines (RBMs) that connect its layers. Starting from the data $\hat{\mathbf{x}}\equiv\hat{\mathbf{s}}^{(0)}$, we train layer $\ell=1,2,\ldots,L$ from the dataset $\hat{\mathbf{s}}^{(\ell-a)}$ by maximising the likelihood
$\mathcal{L}_{\ell}(\theta_{\ell})=\sum_{\mathbf{s}_{\ell-1}}\hat{p}(\mathbf{s}_{\ell-1})\log\sum_{\mathbf{s}_{\ell}}p(\mathbf{s}_{\ell-1},\mathbf{s}_{\ell}|\theta_{\ell})$ (18)
over the parameters $\theta_{\ell}$ of the joint distribution $p(\mathbf{s}_{\ell-1},\mathbf{s}_{\ell}|\theta_{\ell})$ (see Appendix A).
The DBN used for the original datasets is the same as that used in Ref. [37] with ten layers (see Appendix A). Most of our results will focus on layers $\ell=5,6,7,8$ and $9$ and $n_{\ell}=30,25,20,15$ and $10$, respectively. Shallower layers are too high dimensional and their distribution $p(\mathbf{s}_{\ell})$ is hard to estimate numerically.
3.2 Representations approach the HFM with breadth
As a measure of the distance of representations to the HFM we take the Kullback-Leibler (KL) divergence between the empirical distribution of internal layers and the HFM. The empirical distribution is obtained either as the distribution of clamped states, i.e. of states obtained propagating each datapoint through the layers of the DBN, or sampling the distribution by Montecarlo methods.
Notice that there are $2^{n}$ equivalent representations corresponding to gauge transformations $s_{i}\to s_{i}^{\prime}=\tau_{i}s_{i}+(1-\tau_{i})(1-s_{i})$ with $\tau_{i}=0,1$. In order to fix this gauge we set $\mathbf{\tau}$ such that the most frequently sampled state in each layer corresponds to the featureless state $\mathbf{s}_{\ell}^{\prime}=\mathbf{0}$, as for the HFM. In addition, there are $n!$ possible ways to order the variables $\mathbf{s}_{1:n}$. Therefore we find the permutation of the variables for which the KL divergence from the HFM is minimal. The combined effect of these two operations are formally defined by the transformation
$\mathbf{s}^{\prime}=\mathcal{G}_{\tau,\pi}(\mathbf{s}),\qquad s_{i}^{\prime}=\tau_{\pi(i)}s_{\pi(i)}+(1-\tau_{\pi(i)})(1-s_{\pi(i)})$ (19)
where $\pi=(\pi(1),\ldots,\pi(n))$ is a permutation of the integers $1,\ldots,n$. Fig. 3 as well as all other results of this Section are derived performing this transformation.
Fig. 3 plots the KL divergence between the internal representation and the HFM computed on the marginal distributions on the first $n$ variables of both distributions (with $n\leq n_{\ell}$). The reason for this choice is that the KL divergence decreases with the number $n$ of variables for all datasets and hence it decreases with depth. Yet this is a spurious dependence which is due to statistical effects [45]. Marginal distributions make it possible to compare different layers keeping $n$ fixed, thus disentangling the dependence on depth and that on the number of variables.
Figure 3: Fit of the marginal empirical distribution of the first $n$ variables $\mathbf{s}_{1:n}$ of the $\ell^{\rm th}$ layer of a DBN with the marginal distribution $h_{n_{\ell}}(\mathbf{s}_{1:n})$ of the HFM, for $n=2,\ldots,n_{\ell}$ ($n_{\ell}$ is the number of variables of layer $\ell$). Data refer to layers $\ell=5$ to $9$ of a DBN trained on datasets 2-M, M and EM. EMs refers to the sampled distribution of the DBN trained on EM with $10^{6}$ points. All other data refer to clamped distributions. For each dataset, the rightmost datapoint corresponds to the KL divergence of the empirical distribution of the internal layer and the HFM with the same number of variables.
The main message of Fig. 3 is that the broader the data, the more the internal representation approaches the HFM. The 2-M dataset composed of only one digit is clearly further away from the HFM with respect to the dataset M with all ten digits, and the latter is further away from the HFM with respect to the datasets EM with digits and letters. The equilibrium distribution (EMs) appears to approach even more the HFM than the clamped distribution (EM).
Fig. 3 also suggests that the internal representation moves away from the HFM with depth for the narrower dataset (2-M). Such a behaviour would be consistent with the fact that deeper layers extract features which are more and more specific of the digit two. A similar behaviour characterises the deepest layers of the richer datasets, although the statistical evidence is much weaker.
3.3 Multi-peak structure of internal representations
A key difference between the HFM and the empirical distributions $\hat{p}_{\ell}(\mathbf{s}_{\ell})$ of activations of the layers of the DBN is that while the former is composed of a single peak, the latter are characterised by multiple peaks. There are different ways of identifying these peaks. One way is to use TAP equations [46], following Refs. [47]. We leveraged this method, whose details are recalled in Appendix D, to test the theoretical arguments of Section 2.1 suggesting that depth promotes the emergence of peaks in the marginal distributions $p(\mathbf{s}_{\ell})$ learned by the DBN. In order to disentangle the effect of depth from that of width, we trained a DBN with 15 layers on the MNIST dataset, with 250 neurons in each layer. Fig. 4 A) shows that the number of TAP solutions decreases sharply with depth, after the third layer. Fig. 4 B) tests the arguments presented in Section 2.1 showing that for two choices of the function $\phi_{0}(\mathbf{x})$ the distribution of $\phi_{\ell}$ develops sharp peaks with depth.
Figure 4: A) Number of TAP solutions in a DBN with 15 layers of $n=250$ variables, trained on MNIST. B) Joint distribution $p(\psi_{\ell},\phi_{\ell})$ for layers $\ell=6$ and $9$ for the DBN with 10 layers discussed in Appendix A trained on dataset EM. The dependence of the two functions $\psi_{0}(\mathbf{x})$ and $\phi_{0}(\mathbf{x})$ on the coordinates $\mathbf{x}\in[1,28]^{2}$ are shown on the right.
In order to focus on the dominant local maxima, we resort to a different, simpler algorithm. We first sort the configurations $\mathbf{s}_{\ell}$ in descending order of their frequency $\hat{p}_{\ell}(\mathbf{s}_{\ell})$. The topmost configuration $\mathbf{s}^{(0)}={\rm arg}\max_{\mathbf{s}}\hat{p}(\mathbf{s})$ identifies the first peak. Scrolling down the list, states are assigned to the first peak as long as their distance to the closest point belonging to the first peak is smaller than a threshold (that we take to be $n/3$). The first configuration $\mathbf{s}^{(1)}$ whose minimal distance from the first peak exceeds the threshold identifies the top of the second peak. Scrolling further down the list, all configurations are assigned either to the first or the second peak according to the minimal distance to either one. This procedure is iteratively repeated for each of the two peaks as long as the algorithm detects the presence of a second peak. This procedure is illustrated in Fig. 5 for the dataset EMs and it splits the empirical distribution
$\hat{p}(\mathbf{s})=\sum_{\alpha}w_{\alpha}\hat{p}_{\alpha}(\mathbf{s})$ (20)
into a mixture of non-overlapping components $\hat{p}_{\alpha}(\mathbf{s})$. Fig. 5 reports the KL divergence of different peaks from the HFM. It has to be noticed that for each component $\alpha$ a different transformation $\mathcal{G}_{\tau_{\alpha},\pi_{\alpha}}$ is applied to the configurations. This analysis suggests that the peak structure achieves its highest complexity at intermediate layers. This is consistent with the findings of Song et al. [37], who studied the same architecture, and found that layer $\ell=6$ is the one with optimal generative power and ideal resolution-relevance trade-off.
Figure 5: Top left inset: procedure for identifying different peaks in the empirical distribution $\hat{p}(\mathbf{s})$ of DBN’s layers. Main figure: the average KL divergence of the peaks $\hat{p}_{\alpha}(\mathbf{s})$ from the HFM (full line) is plotted against the number $n$ of nodes, for layers $5$ to $9$ of the DBN, for the EMs dataset, from right to left. The KL divergence of each peak from the HFM is also shown, with a colour-code that depends on the entropy $H[\mathbf{s}^{(\ell)}_{\alpha}]$ of $\hat{p}_{\alpha}(\mathbf{s}^{(\ell)}_{\alpha})$. The structure of peaks, as they are revealed by the algorithm, is shown below the curve for the different layers. Layers $8$ and $9$ both have $4$ peaks arranged in a tree of depth two.
The colour code in Fig. 5 reports the values of the entropy $H[\mathbf{s}^{(\ell)}_{\alpha}]$ of different peaks in the various layers. Peaks become sharper (i.e. smaller $H[\mathbf{s}^{(\ell)}_{\alpha}]$) with depth, consistent with the theoretical arguments of the previous Section. Also, for the same layer, peaks with higher $H[\mathbf{s}^{(\ell)}_{\alpha}]$ are closer to the HFM.
Fig. 6 analyzes the first two steps of the procedure leading from one to two peaks, and the next one leading to four peaks. The bottom row of Fig. 6 complements the analysis of the KL divergence with a different distance measure $d(k,m)=1+\tau(k,m)$ based on the Kendall’s $\tau$ correlation between the number of times $k_{\mathbf{s}}$ that a configuration is observed in the sample and $m_{\mathbf{s}}=\max\{i:~{}s_{i}=1\}$. We expect $k_{\mathbf{s}}$ and $m_{\mathbf{s}}$ to be exactly anti-correlated – i.e. $\tau(k,m)=-1$ and hence $d(k,m)=0$ – if $\mathbf{s}$ is sampled from a HFM.
Both measures confirm that internal representations and each of their components approach the HFM as the dataset gets broader (three leftmost panels of Fig. 6). In addition, we find that the decomposition of $\hat{p}(\mathbf{s})$ into components approaches the HFM closer the finer is the decomposition. This shows that individual peaks are closer to the HFM than multi-peak distributions.
Figure 6: Top row: KL divergence between the decomposition of $\hat{p}(\mathbf{s})$ into peaks and the HFM for DBNs trained on datasets 2-M, M and EM, respectively, from left to right, and for an RBMs trained on dataset M with the same number of nodes (rightmost panels). Bottom row: distance $d(k,m)$ for the same data. For 2 (green) and 4 (blue) peaks the solid line reports the average KL divergence from the HFM. For each panel the horizontal axis is the number $n$ of nodes of the layers $5$ to $9$. Note that depth runs from right to left.
In order to probe the effect of depth, Fig. 6 also compares the decomposition of the distribution of different layers of the DBN into a different number of components with the one (rightmost panel of Fig. 6) obtained analysing RBMs with the same number of nodes, trained on the same data (for dataset M). Comparing the rightmost panel of Fig. 6 with the second from the left, we find that the single-peak distribution of the RBM appears closer to the HFM than that of the DBN. Yet for the DBN the distribution of finer components approaches the HFM whereas for RBMs the decomposition has no significant effect on the distance to the HFM. This supports the conjectures that peaks approach the HFM under the combined effect of depth and breadth.
4 Discussion
It has been argued [48] that intelligent behaviour relies on ”extreme generalisation [intended as] the ability to handle entirely new tasks that only share abstract commonalities with previously encountered situations, applicable to any task and domain within a wide scope.” [48]. In this view, the map of “abstract commonalities” that intelligence navigates should necessarily rely on a universal representation, in the limit of an unbounded scope. The HFM is an example of such a universal representation, within the admittedly oversimplified domain of static representations of binary variables. This paper argues that such abstract representations can emerge spontaneously when learning a broader and broader universe of data in deep network architectures. Technically we derive such universal, abstract representation as a fixed point of a RG transformation.
We present numerical experiments corroborating this picture. The range of breadth of the data that these numerical experiments explore is rather limited, yet it is sufficient to confirm that the internal representation of DBNs approach the HFM as the data is drawn from a broader domain. This applies both to the representation as a whole and to the individual “object manifolds” it is composed of, which we identify with the different peaks. This suggest that the distance from the HFM can be taken as a quantitative measure of the level of abstraction of a representation.
Ultimately, the only common characteristic of data coming from very different domains is the coding cost, i.e. the number of bits needed to efficiently code each data point. The principle of maximal relevance predicates that coding costs should be as broadly distributed as possible, which in turns facilitates robust alignment of different data sources along this dimension. The HFM arises as the ideal abstract representation because it is the optimal scaffold for organising data according to their coding cost [^13] Note that the coding cost $-\log h_{n}(\mathbf{s})=gm_{\mathbf{s}}-\log Z_{n}$ depends linearly on the sufficient statistics $m_{\mathbf{s}}$, so the coding cost is itself a sufficient statistics..
Understanding how the conceptual framework developed here can be extended to more complex situations is an interesting avenue of further research [^14] In this respect, we note that the HFM can easily be generalised to variables $x_{i}$ taking value in an arbitrary set $\chi$ by invoking a transformation $\sigma:\chi\to\{0,1\}$ that maps each value of $x_{i}$ into a binary variable $s_{i}=\sigma(x_{i})$ [49]. Extending this analysis in the time dependent domain constitutes a considerably more challenging avenue.. In this vein, the archetypal example of an abstract representation is language. Within the Chomskyan approach to linguistics [50], which has been very influential, one has to distinguish a deep structure – that encodes abstract semantic structures as well as grammatical rules – and a surface structure which is derived from the deep one through a series of transformations leading to the actual, observable form of language as it is spoken or written [50]. The deep structure entails an innate generative process – the universal grammar – which is argued to be common to all human languages, and which relies on the capacity of infinite recursion [51] thus making it possible to generate an infinite variety of sentences with a finite vocabulary. The fact that this capacity emerges in children without exposure to much data (spoken language) [52] has led to the hypothesis that universal grammars need to be biologically hardwired, an hypothesis that is not widely accepted [53]. It is tempting to speculate that universal grammars could emerge in deep cortical areas as fixed points of a transformation such as the one discussed here, driven by the integration of inputs from a broad set of sources, across all sensory modalities. Such universal representations would then be shaped by data which is not limited to language. In this view, it would be the integration of all experience into the same framework – that one may call understanding – that promotes abstraction, with the emergence of universal representations.
5 Acknowledgements
We are grateful to Paolo Muratore and Davide Zoccolan for interesting discussions and to Giulia Betelli for her contribution [49]. We acknowledge Max Planck Research School (IMPRS) for The Mechanisms of Mental Function and Dysfunction (MMFD) for supporting Elias Seiffert.
Appendix A Data, DBNs and their training
A deep belief networks (DBN) consists of Restricted Boltzmann Machines (RBM) stacked one on top of the other, as shown in Fig.7. Each RBM is a Markov random field with pairwise interactions defined on a bipartite graph of two non interacting layers of variables: visible variables $\textbf{x}=(x_{1},..,x_{m})$ representing the data, and hidden variables $\textbf{s}=(s_{1},...,s_{n})$ that are the latent representation of the data. The probability distribution of a single RBM is:
$p(\textbf{x},\textbf{s})=\frac{1}{Z}\exp{\left(\sum_{i,j}W_{ij}x_{i}s_{j}+\sum_{k}x_{k}c_{k}+\sum_{l}s_{l}b_{l}\right)}.$ (21)
where $\textbf{W}=\{W_{ij}\},~{}\textbf{c}=(c_{1},\ldots,c_{m})$ and $\textbf{b}=(b_{1},\ldots,b_{n})$ are the parameters that are learned during training.
Figure 7: A three layer Deep Belief Network
In order to train the DBN we learn the parameters one layer at a time, following the prescription of Hinton [54]. It consists of training the first RBM on the data and then to propagate the input data $\hat{\mathbf{x}}=(\mathbf{x}_{1},\ldots,\mathbf{x}_{N})$ forward to the first hidden layer, thus obtaining a sample of the hidden states $\hat{\mathbf{s}}^{(1)}$ for the first layer. This is then used as input for training the second hidden layer, and so on. This type of training procedure was proven [54] to increase a variational lower bound for the log likelihood of the data set.
In order to generate samples from the trained DBN we consider the connections between the top two layers as undirected, whereas all lower layers are connected to the upper layer by directed connections. This means that, in order to obtain a sample from a DBN we use Gibbs sampling to sample the equilibrium of the top RBM $p_{L}(\textbf{s}^{(L)},\textbf{s}^{(L-1)})$. Then we use this data to sample the states of lower layers using the conditional distribution $p(\mathbf{s}_{\ell-1}|\mathbf{s}_{\ell})$. In this way, we propagate the signal till the visible layer.
The DBN used in our experiment is the same as that used in Ref. [37]: it has a visible layer with $784$ nodes and $L=10$ hidden layers with the following number of nodes: $n_{\ell}=500,250,120,60,30,25,20,15,10$ and $5$, for $\ell=1,\ldots,10$. The MNIST [55], eMNIST [56] The datasets were accessed via the torchvision library [57].
In order to learn the parameters of a single RBM we used stochastic gradient ascent on the log-likelihood, employing Persistent Contrastive Divergence with $k=10$, a learning rate of $0.01$, and mini-batches of size $64$ (see [58]), for $\sim 10^{5}$ epochs. After the training, the clamped states $\{\hat{\mathbf{s}}^{\ell}\}_{\ell=1}^{L}$ was obtained starting from the input data $\hat{\mathbf{x}}$ at the first layer and sequentially sampling each subsequent layer from its corresponding conditional distribution. In contrast, the equilibrium sample $\hat{\mathbf{s}}^{\ell}$ of a given layer $\ell$ was obtained by sampling the equilibrium distribution of $p_{\ell}(\mathbf{s}_{\ell},\mathbf{s}_{\ell-1})$. This was achieved by initializing with a random configuration $\hat{\mathbf{s}}^{\ell}_{0}$, and performing alternating Gibbs sampling for $10^{6}$ steps to ensure convergence to the equilibrium distribution.
Decelle et al. [59] [60] have shown that the distribution learned by an RBM trained with CD-10 does not reproduce equilibrium distribution, but it can still serve as a good generative model when sampled out of equilibrium. Instead they observed that persistent contrastive divergence (PCD-10), the algorithm we used, learns a good equilibrium distribution [^15] In Contrastive Divergence-k (CD-k), the Markov chain used to sample the distribution is initialized on the batch used to compute the gradient and $k$ Monte Carlo steps are performed. In Persistent Contrastive Divergence-k (PCD-k) the MCMC is initialized in the configuration of the previous epoch.
Appendix B Existence and uniqueness of the solution for $\alpha$
Each step in the coarse graining RG involves a change in entropy $\Delta_{k\to k+1}H=H_{k+1}[s]-H_{k}[s]$ as follows:
$\Delta_{1\to 2}H$ $=$ $H[s_{1:n-1}]-H[s_{1:n}]=-H[s_{n}|s_{1:n-1}]$ (22)
[…]
$\Delta_{3\to 4}H$ $=$ $h(q)-\alpha H[s_{1:n}]-(1-\alpha)h(\tilde{p}_{n}(0_{1:n})),\qquad q=\alpha+(1-\alpha)\tilde{p}_{n}(0_{1:n})$ (24)
where $h(x)=-x\log_{2}x-(1-x)\log_{2}(1-x)$. Overall the change in entropy is
$\Delta_{1\to 4}H=h(q)-\alpha H[s_{1:n}]-(1-\alpha)h(\tilde{p}_{n}(0_{1:n}))+1-H[s_{n}|s_{1:n-1}]$ (25)
therefore $\alpha$ is the solution of the equation $\Delta_{1\to 4}H=0$. Notice that $\Delta_{1\to 4}H(\alpha=0)\geq 0$ and
$\frac{d}{d\alpha}\Delta_{1\to 4}H=(1-\tilde{p}_{n}(0_{1:n}))\log\frac{1-q}{q}-H[s_{n}|s_{1:n-1}]+h(\tilde{p}_{n}(0_{1:n}))$ (26)
which is negative at $\alpha=0$ and
$\frac{d^{2}}{d\alpha^{2}}\Delta_{1\to 4}H=-\frac{1-\tilde{p}_{n}(0_{1:n})}{q(1-q)}<0$ (27)
which means that the solution it is unique provided that $\Delta_{1\to 4}H(\alpha=1)=-H[\mathbf{s}_{1:n}]+1-H[s_{n}|\mathbf{s}_{1:n-1}]$ is negative. A sufficient condition is that $H[\mathbf{s}_{1:n}]>1$.
Appendix C Proof of Eq. (12)
Let us first analyse how the HFM transforms under $\Re_{\uparrow}$. Marginalisation on $s_{n}$ yields
$h_{n}(\mathbf{s}_{1:n-1})=\sum_{s_{n}=0,1}h_{n}(\mathbf{s}_{1:n})=\frac{Z_{n-1}}{Z_{n}}h_{n-1}(\mathbf{s}_{1:n-1})+\frac{1}{Z_{n}}e^{-gn}$ (28)
Hence
$\tilde{p}_{n}(\mathbf{s}_{1:n}^{\prime})=\frac{Z_{n-1}}{2Z_{n}}h_{n-1}(\mathbf{s}_{2:n}^{\prime})+\frac{1}{2Z_{n}}e^{-gn}$ (29)
If $\mathbf{s}_{2:n}^{\prime}=0$ then
$h_{n-1}(\mathbf{s}_{2:n}^{\prime})=\frac{Z_{n}}{Z_{n-1}}h_{n}(\mathbf{s}_{1:n}^{\prime})e^{gs_{1}^{\prime}}=\frac{Z_{n}}{Z_{n-1}}h_{n}(\mathbf{s}_{1:n}^{\prime})\left[e^{g}+(1-e^{g})\delta_{s_{1}^{\prime},0}\right]$
because $m_{\mathbf{s}_{1:n}}=s_{1}$ in this case. If $\mathbf{s}_{2:n}^{\prime}\neq\mathbf{0}_{2:n}$ instead
$h_{n-1}(\mathbf{s}_{2:n}^{\prime})=\frac{Z_{n}}{Z_{n-1}}e^{g}h_{n}(\mathbf{s}_{1:n}^{\prime})$
because $m_{\mathbf{s}_{2:n}}=m_{\mathbf{s}_{1:n}}-1$. Therefore, both cases are accounted for by the equation
$h_{n-1}(\mathbf{s}_{2:n}^{\prime})=\frac{Z_{n}}{Z_{n-1}}e^{g}h_{n}(\mathbf{s}_{1:n}^{\prime})+\frac{1-e^{g}}{Z_{n-1}}\delta_{\mathbf{s}_{1:n}^{\prime},0}$ (30)
Substituting this into Eq. (29) yields
$\tilde{p}(\mathbf{s}_{1:n}^{\prime})=\frac{e^{g}}{2}h_{n}(\mathbf{s}_{1:n}^{\prime})+\frac{1-e^{g}}{2Z_{n}}\delta_{\mathbf{s}_{1:n}^{\prime},0}+\frac{e^{-gn}}{2Z_{n}}$ (31)
and
$p_{n}^{\prime}(\mathbf{s}_{1:n}^{\prime})=(1-\alpha)\frac{e^{g}}{2}h_{n}(\mathbf{s}_{1:n}^{\prime})+\left[\alpha-(1-\alpha)\frac{e^{g}-1}{2Z_{n}}\right]\delta_{\mathbf{s}_{1:n}^{\prime},0}+(1-\alpha)\frac{e^{-gn}}{2Z_{n}}$ (32)
Therefore, at least for finite $n$, the HFM is not a fixed point.
We look for a fixed point of the form
$p_{n}^{*}(s_{1:n})=(1-\beta)h_{n}(s_{1:n})+\beta u_{n}(s_{1:n})$ (33)
exploiting the fact that $\Re_{\uparrow}$ is a linear transformation. The uniform distribution $u_{n}(s_{1:n})=2^{-n}$ transforms as $\Re_{\uparrow}(u_{n})(\mathbf{s}_{1:n})=(1-\alpha)2^{-n}+\alpha\delta_{\mathbf{s}_{1:n},0}$.
After some calculation, with $\xi=2e^{-g}$, we find
$p_{n}^{\prime}(\mathbf{s}_{1:n}^{\prime})$ $=$ $\frac{(1-\beta)(1-\alpha)}{\xi}h_{n}(\mathbf{s}_{1:n}^{\prime})$ (34)
$+\left[(1-\beta)\left(\alpha-(1-\alpha)\frac{2-\xi}{2\xi Z_{n}}\right)+\beta\alpha\right]\delta_{\mathbf{s}_{1:n}^{\prime},0}$ (35)
$+\left[\frac{(1-\beta)(1-\alpha)}{2Z_{n}}{\xi^{n}}+\beta(1-\alpha)\right]u_{n}(\mathbf{s}_{1:n}^{\prime})$ (36)
Setting the coefficient of $h_{n}(s_{1:n}^{\prime})$ in the first line (34) to $1-\beta$ and the coefficient of $u_{n}(s_{1:n}^{\prime})$ in the third line  (36) equal to $\beta$ yields
$\alpha=1-\xi,\qquad\beta=\frac{\xi^{n+1}}{2-\xi}$ (37)
the second line (35) then vanishes by normalization. The solution then reads
$p^{*}_{n}(\mathbf{s}_{1:n})=\left(1-\frac{1}{e^{g}-1}\right)e^{-gm_{\mathbf{s}_{1:n}}}+\frac{1}{e^{g}-1}e^{-gn}\,.$ (38)
Interestingly, a solution only exists for $\xi<1$, i.e. for $g>g_{c}$, and in the limit $g\to g_{c}$ the fixed point distribution tends to $u_{n}$. Eq. (38) has the same form of an HFM with $m>n$ features, marginalised over the $m-n$ most detailed ones (see Eq. 5). The value of $m$ can be computed equating $1-\beta$ to the coefficient of $u_{n}$ in the marginal of $h_{m}(\mathbf{s}_{1:m})$ over $\mathbf{s}_{1:n}$. This yields the equation
$\frac{\xi^{m+1}(2-\xi-\xi^{n+1})}{(2-\xi)(2-\xi-\xi^{m+1})}=0$ (39)
whose only solution for $\xi<1$ is $m=+\infty$. In other words, the fixed point $p_{n}^{*}$ is the marginal distribution of the $n$ most coarse grained features of an HFM with infinite features, which is Eq. (12).
Appendix D TAP solutions
The Thouless-Anderson-Palmer (TAP) equations [46] are the local minima of the TAP free energy, that can be obtained with a high temperature expansion of the Legendre transform of the free energy of an extended system with extra fields on each spin variable. For a RBM the extended free energy takes the form [47]:
$-\beta F(\boldsymbol{\phi},\boldsymbol{\psi})=\log\sum_{\mathbf{x,s}}\exp\bigg{[}-\beta E(\mathbf{x,s;W,b,c})+\sum_{i}\phi_{i}x_{i}+\sum_{\mu}\psi_{\mu}s_{\mu}\bigg{]}\,.$ (40)
The Legendre transform rephrases the problem in the space of magnetizations:
$-\beta\Gamma_{\beta}(\mathbf{m^{x},m^{s}})=-\beta\max_{\boldsymbol{\phi},\boldsymbol{\psi}}\bigg{[}F(\boldsymbol{\phi},\boldsymbol{\psi})+\sum_{i}\phi_{i}m_{i}^{x}+\sum_{\mu}\psi_{\mu}m^{s}_{\mu}\bigg{]}$ (41)
The TAP equations for the RBM magnetizations can be obtained from the stationary conditions of a second order expansion of $\Gamma_{\beta}(\mathbf{m^{x},m^{s}})$ around $\beta=0$, as was done in [47]:
$m_{j}^{s}$ $=\sigma\bigg{[}b_{j}+\sum_{i}W_{ij}m_{i}^{x}-W_{ij}^{2}\bigg{(}m_{j}^{s}-\frac{1}{2}\bigg{)}(m_{i}^{x}-(m_{i}^{x})^{2})\bigg{]}$
$m_{i}^{x}$ $=\sigma\bigg{[}c_{i}+\sum_{j}W_{ij}m_{j}^{s}-W_{ij}^{2}\bigg{(}m_{i}^{x}-\frac{1}{2}\bigg{)}(m_{j}^{s}-(m_{j}^{s})^{2})\bigg{]}$ (42)
Solutions to these equations for a given layer can be found by using an iterative algorithm. We use as initialization the clamped activations of each layer for each datapoint in the MNIST datatset.